# Mora: Enabling Generalist Video Generation via A Multi-Agent Framework

### v2

@@ 4.2 Results
Text-to-Video Generation.  The quantitative results are detailed in Table 2. Mora showcases commendable performance across all metrics, making it highly comparable to the top-performing model, Sora, and surpassing the capabilities of other competitors. Specifically, Mora achieved a Video Quality score of 0.792, which closely follows Sora’s leading score of 0.797 and surpasses the current best open-source model like VideoCrafter1. In terms of Object Consistency, Mora scored 0.95, equaling Sora and demonstrating superior consistency in maintaining object identities throughout the videos. For Background Consistency and Motion Smoothness, Mora achieved scores of 0.95 and 0.99, respectively, indicating high fidelity in background stability and fluidity of motion within generated videos. Although Sora achieved 0.96 slightly outperforms Mora in Background Consistency, the margin is minimal. The Aesthetic Quality metric, which assesses the overall visual appeal of the videos, saw Mora scoring 0.57. This score, while not the highest, reflects a competitive stance against other models, with Sora scoring slightly higher at 0.60. Nevertheless, Mora’s performance in Dynamic Degree and Imaging Quality, with scores of 0.70 and 0.59, showcases its strength in generating dynamic, visually compelling content that surpasses all other models. As for Temporal Style, Mora scored 0.26, indicating its robust capability in addressing the temporal aspects of video generation. Although this performance signifies a commendable proficiency, it also highlights a considerable gap between our model and Sora, the leader in this category with a score of 0.35.
@@
Text-conditional Image-to-Video Generation.  Analyzing Table 3, we discern a notable demonstration of Mora’s capabilities in text-conditional image-to-video generation, closely trailing behind Sora. Sora leads with a VideoTI score of 0.90 and Motion Smoothness of 0.99, underscoring its refined alignment of video with the provided text and the fluidity of motion within its generated videos. Mora’s VideoTI score of 0.88 and Motion Smoothness of 0.97, while marginally lower, reflect its robust potential in the accurate interpretation of textual prompts and the generation of smooth motion sequences. Both models exhibit a matched score in Dynamic Degree at 0.75, indicating a comparable proficiency in animating videos with a sense of activity and movement. The Imaging Quality scores show Sora with a slight lead at 0.63 against Mora’s 0.60, suggesting Sora’s subtle superiority in image rendering within video sequences.
Table 3: Text-condtional image-to-video generation. This task evaluates the model’s ability to convert static images and accompanying textual instructions into coherent video sequences. We used 4 pairs of image and text inputs sourced from the official Sora technical report OpenAI (2024b).
Model VideoTI Motion Smoothness Dynamic Dgree Imaging Quality
Sora 0.90 0.99 0.75 0.63
Mora 0.88 0.97 0.75 0.60
@@
In Figure 3, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image. The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models. Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video. This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
Extend Generated Videos.  The quantitative results from Table 4 reveal that while Sora holds a slight edge over Mora in TCON and Imaging Quality, indicating a higher consistency and fidelity in extending video sequences, Mora’s performance remains close, with a marginal difference of 0.05 in TCON and 0.04 in Imaging Quality. The closeness in Temporal Style scores, 0.24 for Sora and 0.22 for Mora, further signifies that Mora nearly matches Sora in maintaining stylistic continuity over time. Despite Sora’s lead, Mora’s capabilities, particularly in following the temporal style and extending existing videos without significant quality loss, demonstrate its effectiveness in the video extension domain.
Table 4: Extend generated videos. This experiment compares the continuity and quality of the extended video with the same input video given to Sora and Mora respectively. The input video is taken from the full video of the offical Sora technical report OpenAI (2024b), and we take 2 clips as the input video for this task.
Model TCON Imaging Quality Temporal Style
Sora 0.99 0.43 0.24
Mora 0.94 0.39 0.22
@@
From a qualitative standpoint, Figure 4 illustrates the competencies of Mora in extending video sequences. Both Sora and Mora adeptly maintain the narrative flow and visual continuity from the original to the extended video. Despite the slight numerical differences highlighted in the quantitative analysis, the qualitative outputs suggest that Mora’s extended videos preserve the essence of the original content with high fidelity. The preservation of dynamic elements such as the rider’s motion and the surrounding environment’s blur effect in the Mora generated sequences showcases its capacity to produce extended videos that are not only coherent but also retain the original’s motion and energy characteristics. This visual assessment underscores Mora’s proficiency in generating extended video content that closely mirrors the original, maintaining the narrative context and visual integrity, thus providing near parity with Sora’s performance.
Table 5: Video-to-video editing. This task examines the model’s ability to edit and transform existing video content in accordance with textual instructions while maintaining visual and stylistic coherence. Data includes 1 input video and 12 instructions from official Sora technical report.
Model Imaging Quality Temporal Style
Sora 0.52 0.24
Mora 0.38 0.23
Video-to-Video Editing.  Table 5 exhibits a comparative analysis of video-to-video editing capabilities between Sora and Mora. Sora secures a higher score in Imaging Quality at 0.52, which suggests a superior capability in preserving the visual details and overall image fidelity during the video editing process. Mora demonstrates capability in video-to-video editing with an Imaging Quality score of 0.38. Although this score reveals a noticeable discrepancy when compare to Sora, it offers valuable insights on pinpointing areas for targeted enhancements in future iterations of the Mora. In the aspect of Temporal Style, both models exhibit proximal performance, with Sora marginally leading at 0.24 compared to Mora’s 0.23. This near-parity underscores the capacity of Mora to closely emulate the stylistic temporal consistency achieved by Sora, ensuring that the edited videos retain a coherent style over time, a crucial aspect of seamless video-to-video editing.
@@
Table 6: Connect Videos. This task assess the capability of models to seamlessly integrate distinct video clips into a cohesive sequence. $*{}^{*}$ora data we utilize is 5 distinct video-video pairs from Sora technical report website and Mora data we utilize 5 video-video pairs of similar styles.
Model Imaging Quality Tmean
$*{}^{*}$ 0.52 0.64
Mora $*{}^{*}$ 0.42 0.45
Figure 6: Samples for Connect Videos
Upon qualitative evaluation, Figure 5 presents samples from video-to-video editing tasks, wherein both Sora and Mora were instructed to modify the setting to the 1920s style while maintaining the car’s red color. Visually, Sora’s output exhibits a transformation that convincingly alters the modern-day setting into one reminiscent of the 1920s, while carefully preserving the red color of the car. Mora’s transformation, while achieving the task instruction, reveals differences in the execution of the environmental modification, with the sampled frame from generated video suggesting a potential for further enhancement to achieve the visual authenticity displayed by Sora. Nevertheless, Mora ’s adherence to the specified red color of the car underline its ability to follow detailed instructions and enact considerable changes in the video content. This capability, although not as refined as Sora’s, demonstrates Mora’s potential for significant video editing tasks.
Connect Videos.  Quantitative analysis based on the results shown in Table 6 suggests that the model Sora outperforms Mora in terms of Imaging Quality and Tmean. Sora achieves a score of 0.52 in Imaging Quality, indicating its higher fidelity in visual representation compared to Mora’s 0.42. Furthermore, Sora’s superiority is evident in the Temporal coherence aspect with a score of 0.64, which implies that Sora maintains a more consistent visual narrative over time than Mora, which scores 0.45. This quantitative assessment not only solidifies Sora’s position as a superior model in creating high-quality, temporally coherent video sequences but also delineates a trajectory for future enhancements in video connectivity for the Mora framework.
Qualitative analysis based on Figure 6 suggest that, in comparison to Sora’s proficiency in synthesizing intermediate video segments that successfully incorporate background elements from preceding footage and distinct objects from subsequent frames within a single frame, the Mora model demonstrates a blurred background in the intermediate videos, which results in indistinguishable object recognition. Accordingly, this emphasizes the potential for advancing the fidelity of images within the generated intermediate videos as well as enhancing the consistency with the entire video sequence. This would contribute to refining the video connecting process and improving the integration quality of Mora’s model outputs.
Table 7: Simulate digital worlds. This task assesses model’s effectiveness in creating videos that emulate digital or virtual environments, with a focus on preserving the visual details and the distinct appearance style of such worlds. By referring to the video generated by Sora from the official technical report OpenAI (2024b), we used GPT-4 OpenAI (2023b) to build the input prompt for Mora, which contains the "minecraft scene" and the "player character".
Model Imaging Quality Appearance style
Sora 0.62 0.23
Mora 0.52 0.23
@@
Simulate Digital Worlds.  In the assessment of digital world simulation capabilities, Table 7 presents a comparative metric analysis between Sora and Mora. Sora exhibits a lead in Imaging Quality with a score of 0.62, indicative of its refined capability to render digital worlds with a higher degree of visual realism and fidelity. Mora, with a score of 0.52, although demonstrating a competent level of performance, falls behind Sora, suggesting areas for improvement in achieving the same level of image clarity and detail. However, both models achieve identical scores in Appearance Style, at 0.23, which reflects a shared ability to adhere to the stylistic parameters of the digital worlds being simulated. This suggests that while there is a difference in the imaging quality, the stylistic translation of textual descriptions into visual aesthetics is accomplished with equivalent proficiency by both models.
@@
Upon qualitative evaluation, Figure 7 presents samples from Simulate digital worlds tasks, wherein both Sora and Mora were instructed to generated video of "Minecraft" scenes. In the top row of frames generated by Sora, we see that the videos maintain high fidelity to the textures and elements typical of digital world aesthetics, characterized by crisp edges, vibrant colors, and clear object definition. The pig and the surrounding environment appear to adhere closely to the style one would expect from a high-resolution game or a digital simulation. These are crucial aspects of performance for Sora, indicating a high-quality synthesis that aligns well with user input while preserving visual consistency and digital authenticity. The bottom row of frames generated by Mora suggests a step towards achieving the digital simulation quality of Sora but with notable differences. Although Mora seems to emulate the digital world’s theme effectively, there is a visible gap in visual fidelity. The images generated by Mora exhibit a slightly muted color palette, less distinct object edges, and a seemingly lower resolution compared to Sora’s output. This suggests that Mora is still in a developmental phase, with its generative capabilities requiring further refinement to reach the performance level of Sora.
5 Discussion
5.1 Strengths of Mora
Innovative Framework and Flexibility.  Mora introduces a groundbreaking multi-agent framework for video generation, significantly advancing the field by enabling a vast array of tasks. This innovative approach not only facilitates text-to-video conversion but also supports the simulation of digital worlds, showcasing unparalleled flexibility and efficiency. Unlike its closed-source counterparts such as Sora, Mora’s open framework design stands out, offering the ability to seamlessly integrate various models. This adaptability allows for the completion of an expanded range of tasks and projects, making Mora an indispensable tool for diverse application scenarios.
Open-Source Contribution.  Mora’s open-source nature is highlighted as a significant contribution to the AI community, encouraging further development and refinement by providing a solid foundation upon which future research can build. This open-source approach not only democratizes access to advanced video generation technologies but also fosters collaboration and innovation in the field. The discussion suggests new ways for future research, including improving the framework’s efficiency, reducing its computational demands during training, and exploring new agent configurations to enhance performance.
5.2 Limitations of Mora
Video Dataset is all you need.  Collecting high-quality video datasets poses significant challenges, primarily due to copyright restrictions on many videos. Unlike images, which can often be easier to gather and utilize for training purposes due to a broader range of available resources and more lenient copyright laws, videos frequently contain copyrighted materials that are not as straightforward to collect or use legally. This issue is compounded by the fact that high-quality videos often come from professional sources, such as movies, television shows, and proprietary game footage, which are rigorously protected by copyright laws. Especially in scenarios involving humans in realistic settings, Mora struggles to generate lifelike movements, such as walking or riding a bicycle. Capturing the subtleties of human motion requires not just any video footage, but high-resolution, smoothly captured sequences that detail every aspect of human kinetics, including the nuances of balance, posture, and interaction with surroundings. Without access to extensive datasets that accurately represent this wide range of human movements, it becomes challenging for video generation models like Mora to replicate these actions convincingly. This limitation underscores the importance of not only the quantity but also the quality and diversity of video datasets in training models to understand and recreate complex human behaviors accurately.
Quality and Length Gaps.  Despite its innovative approach, Mora faces notable challenges, especially when compared to its counterpart, Sora, in terms of video generation quality and capabilities. While Mora is capable of accomplishing tasks similar to those of Sora, the quality of videos generated by Mora falls significantly short, particularly in scenarios involving substantial object movement. Such conditions often introduce a considerable amount of noise, with the quality degradation becoming more pronounced relative to the video length. Moreover, although Mora generally maintains quality for up to 12 seconds of video, any attempt to extend beyond this duration leads to a marked decline in video quality. In contrast, Sora has demonstrated the ability to produce high-quality videos exceeding one minute in length, as showcased in its technical report. This disparity in imaging quality, alongside the limitations in generating videos longer than 12 seconds at a resolution of 1024x576, underscores the urgent need for advancements in Mora’s rendering capabilities and an extension of its video generation parameters.
Instruction Following Capability.  Mora has reached a milestone in its development, now capable of generating videos exceeding 10 seconds from given prompts. However, despite its ability to include all objects specified in the prompts within the generated videos, Mora encounters limitations in executing certain functions. Notably, it struggles with interpreting and rendering the dynamics of motion described in prompts, such as the speed of movement. Additionally, Mora lacks the capability to control the direction of motion—specific directions like left or right for specified objects remain unachievable. This shortfall primarily stems from the system’s foundational approach to video generation, which operates on an Image to Video basis without direct input from textual prompts. Consequently, meeting these specific requirements poses significant challenges, highlighting areas for potential enhancement in Mora’s functionality and interpretative capabilities.
Human Visual Preference Alignment.  The absence of human labeling information within the video domain suggests that experimental results may not always align with human visual preferences, highlighting a significant gap. For example, during connection video task, the generation of transition videos that disintegrate a male figure only to form a female figure represents an exceedingly illogical scenario. This example underscores the necessity for datasets that adhere more closely to physical laws, thereby refining our work and ensuring more realistic and coherent video generation outcomes. Such a framework would enhance the appeal of generated videos to viewers and contribute to establishing comprehensive benchmarks and evaluation criteria for the field.
6 Conclusion
We introduce Mora, a groundbreaking generalist framework for video generation that addresses a range of video-related tasks. Leveraging the collaborative power of multiple agents, Mora marks a considerable advancement in generating videos from textual prompts, establishing new benchmarks for adaptability, efficiency, and output quality in the field of video generation. Our thorough evaluation reveals that Mora not only competes with but also exceeds the capabilities of current leading models in certain areas. However, it has a notable gap with Sora model by OpenAI whose closed-source nature poses considerable challenges for replication and innovation within the academic and professional communities. Our work demonstrates the untapped potential of a meta-programming approach that facilitates intricate collaborations among a variety of agents, each specializing in a segment of the video generation process. In essence, the achievements of Mora not only illustrate the current open source state-of-the-art models in video generation but also illuminate the path forward for the field. As we continue to explore the vast landscape of generative AI, multi-agent collaboration frameworks like Mora will undoubtedly play a pivotal role in unlocking new creative possibilities and applications, from storytelling and content creation to simulation and training. The journey of innovation is far from over, and Mora represents a significant milestone on this ongoing voyage of discovery.
Looking ahead, there are several promising pathways for further research. One such direction involves exploring the integration of more sophisticated natural language understanding capabilities within the agents, potentially allowing for more detailed and context-aware video generations. In addition, the expansion of Mora to incorporate real-time feedback loops could offer interactive video creation experiences, where user inputs could guide the generation process in more dynamic and responsive ways. Furthermore, the challenge of accessibility and computational resource requirements remains a critical barrier to wider adoption and innovation. Future iterations of Mora could benefit from optimizations that reduce these requirements, making advanced video generation technologies more accessible to a broader range of users and developers. In parallel, efforts to create more open and collaborative research environments could accelerate progress in this domain, enabling the community to build upon the foundation laid by Mora framework and other pioneering works.

### v3

@@ B.1 More Results Details
Text-to-Video Generation.  The quantitative results are detailed in Table 1 and Figure 7. Mora showcases commendable performance across all metrics, making it highly comparable to the top-performing model, Sora, and surpassing the capabilities of other competitors. Specifically, Mora achieved a Video Quality score of 0.792, which closely follows Sora’s leading score of 0.797 and surpasses the current best open-source model like VideoCrafter1. In terms of Object Consistency, Mora scored 0.95, equaling Sora and demonstrating superior consistency in maintaining object identities throughout the videos. For Background Consistency and Motion Smoothness, Mora achieved scores of 0.95 and 0.99, respectively, indicating high fidelity in background stability and fluidity of motion within generated videos. Although Sora achieved 0.96 slightly outperforms Mora in Background Consistency, the margin is minimal. The Aesthetic Quality metric, which assesses the overall visual appeal of the videos, saw Mora scoring 0.57. This score, while not the highest, reflects a competitive stance against other models, with Sora scoring slightly higher at 0.60. Nevertheless, Mora’s performance in Dynamic Degree and Imaging Quality, with scores of 0.70 and 0.59, showcases its strength in generating dynamic, visually compelling content that surpasses all other models. As for Temporal Style, Mora scored 0.26, indicating its robust capability in addressing the temporal aspects of video generation. Although this performance signifies a commendable proficiency, it also highlights a considerable gap between our model and Sora, the leader in this category with a score of 0.35.
The results in Table 4 compare various models on video generation performance using the Vbench dataset. Mora (Open-Sora-Plan) achieves the highest overall video quality (0.848) and excels in both object and background consistency (0.98), while also leading in aesthetic quality (0.70) and imaging quality (0.72). Emu3 and Gen-3-Alpha perform similarly in video quality (0.841), with Emu3 showing superior motion smoothness (0.99) and dynamic degree (0.79). Despite a slightly lower video quality, LaVie-2 outperforms other models in temporal style and maintains strong consistency metrics across the board.
@@
Figure 8: Demo of Mora on various tasks.
Mora can deal with various tasks, including video editing, video extension, and stimulate digital world. We provide demo of Mora in Figure 8.
Text-conditional Image-to-Video Generation.   In Figure 14, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image. The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models. Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video. This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
@@
Table 4: Comparative analysis of text-to-video generation performance for models with Vbench data.
Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Degree Imaging Quality Temporal Style Video Length(s)
Emu3 0.841 0.95 0.98 0.99 0.60 0.79 0.63 0.24 5
Open-Sora-Plan_V1.2 0.814 0.97 0.98 0.99 0.59 0.42 0.57 0.25 8
Gen-3-Alpha 0.841 0.97 0.97 0.99 0.63 0.60 0.67 0.25 10
LaVie-2 0.832 0.98 0.98 0.98 0.68 0.31 0.70 0.25 3
Mora (Open-Sora-Plan) 0.848 0.98 0.98 0.99 0.70 0.72 0.72 0.25 12
Extend Generated Videos.  From a qualitative standpoint, Figure 10 illustrates the competencies of Mora in extending video sequences. Both Sora and Mora adeptly maintain the narrative flow and visual continuity from the original to the extended video. Despite the slight numerical differences highlighted in the quantitative analysis, the qualitative outputs suggest that Mora’s extended videos preserve the essence of the original content with high fidelity. The preservation of dynamic elements such as the rider’s motion and the surrounding environment’s blur effect in the Mora generated sequences showcases its capacity to produce extended videos that are not only coherent but also retain the original’s motion and energy characteristics. This visual assessment underscores Mora’s proficiency in generating extended video content that closely mirrors the original, maintaining the narrative context and visual integrity, thus providing near parity with Sora’s performance.
@@
Video-to-Video Editing.  Upon qualitative evaluation, Figure 11 presents samples from video-to-video editing tasks, wherein both Sora and Mora were instructed to modify the setting to the 1920s style while maintaining the car’s red color. Visually, Sora’s output exhibits a transformation that convincingly alters the modern-day setting into one reminiscent of the 1920s, while carefully preserving the red color of the car. Mora’s transformation, while achieving the task instruction, reveals differences in the execution of the environmental modification, with the sampled frame from generated video suggesting a potential for further enhancement to achieve the visual authenticity displayed by Sora. Nevertheless, Mora ’s adherence to the specified red color of the car underline its ability to follow detailed instructions and enact considerable changes in the video content. This capability, although not as refined as Sora’s, demonstrates Mora’s potential for significant video editing tasks.
@@
Figure 12: Samples for Video Connetion
Connect Videos.   Qualitative analysis based on Figure 12 suggest that, in comparison to Sora’s proficiency in synthesizing intermediate video segments that successfully incorporate background elements from preceding footage and distinct objects from subsequent frames within a single frame, the Mora model demonstrates a blurred background in the intermediate videos, which results in indistinguishable object recognition. Accordingly, this emphasizes the potential for advancing the fidelity of images within the generated intermediate videos as well as enhancing the consistency with the entire video sequence. This would contribute to refining the video connecting process and improving the integration quality of Mora’s model outputs.
@@
Simulate Digital Worlds.   Upon qualitative evaluation, Figure 13 presents samples from Simulate digital worlds tasks, wherein both Sora and Mora were instructed to generated video of "Minecraft" scenes. In the top row of frames generated by Sora, we see that the videos maintain high fidelity to the textures and elements typical of digital world aesthetics, characterized by crisp edges, vibrant colors, and clear object definition. The pig and the surrounding environment appear to adhere closely to the style one would expect from a high-resolution game or a digital simulation. These are crucial aspects of performance for Sora, indicating a high-quality synthesis that aligns well with user input while preserving visual consistency and digital authenticity. The bottom row of frames generated by Mora suggests a step towards achieving the digital simulation quality of Sora but with notable differences. Although Mora seems to emulate the digital world’s theme effectively, there is a visible gap in visual fidelity. The images generated by Mora exhibit a slightly muted color palette, less distinct object edges, and a seemingly lower resolution compared to Sora’s output. This suggests that Mora is still in a developmental phase, with its generative capabilities requiring further refinement to reach the performance level of Sora.
[…]
Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Degree Imaging Quality Temporal Style
Mora (Open-Sora-Plan) 0.800 0.98 0.97 0.99 0.66 0.50 0.70 0.31
Mora (Open-Sora-Plan) 0.767 0.94 0.95 0.99 0.61 0.43 0.68 0.26
Mora w/o Human-in-the-loop 0.785 0.98 0.96 0.95 0.63 0.50 0.69 0.31
Mora w/o Self-modulated 0.776 0.96 0.95 0.95 0.62 0.51 0.67 0.27
Mora with RI-modulated 0.797 0.98 0.96 0.99 0.66 0.50 0.69 0.29
B.2 Ablation Study
The results in Table 5 clearly demonstrate the effectiveness of each component in our Mora model. The full version of Mora (Open-Sora-Plan) outperforms all ablated variants across most metrics, showcasing the importance of each design choice. First, Mora achieves the highest video quality score of 0.800, along with the best object consistency (0.98), background consistency (0.97), motion smoothness (0.99), and imaging quality (0.70). This underscores the significance of incorporating all components, as removing any one of them leads to a noticeable drop in performance. For instance, removing the human-in-the-loop module (Mora w/o Human-in-the-loop) reduces video quality to 0.785 and slightly decreases motion smoothness (0.95), indicating that human guidance is key to maintaining higher video generation quality and consistency. Similarly, removing self-modulation (Mora w/o Self-modulated) results in a further decline in video quality to 0.776, while slightly improving the dynamic degree (0.51). However, this increase in dynamic degree comes at the cost of consistency across other metrics, showing that self-modulation balances aesthetic quality and consistency. The Random Initial modulation (Mora with RI-modulated) also demonstrates a strong performance, achieving 0.797 in video quality and maintaining the same level of object consistency (0.98) and motion smoothness (0.99) as the full model. However, the overall consistency and quality remain lower than the full Mora model, confirming that learned modulation embeddings contribute significantly to the model’s performance. These results demonstrate that each component—human-in-the-loop, self-modulation, and optimized modulation embeddings—plays a crucial role in improving the model’s text-to-video generation capabilities. The full Mora model’s superior performance validates the importance of our design, providing a clear advantage over the ablated versions.
Table 6: Framework comparison capabilities of Sora, Mora and other autonomous agent framework. "✓“ indicates the specific feature in the corresponding framework or model. "-“ means absence.
Framework Role-based agent SOPs Human in the loop Task
AutoGPT (Aut, ) - - - Code Generation
AutoGen (Wu et al., 2023b) - Code Generation
MetaGPT (Hong et al., 2023) - Code Generation
Sora - - - Video Generation
Mora Video Generation
Appendix C Capabilities Analysis
Compared to the closed-source baseline model Sora and autonomous agents such as MetaGPT (Hong et al., 2023), our framework, Mora, offers enhanced functionalities for video generation tasks. As shown in Table 6, Mora encompasses a comprehensive range of capabilities designed to handle diverse and specialized video generation tasks effectively. The integration of Standard Operating Procedures (SOPs) such as role-play expertise, structured communication, and streamlined workflows, along with human-in-the-loop systems, significantly refines the control and quality of video generation. While other baseline methods can adapt SOP-like designs to boost their performance, they typically do so only within the realm of code generation tasks. In contrast, models like Sora lack the capability for fine-grained control or autonomous video generation, highlighting the advanced capabilities of Mora in this domain.
A detailed comparison of tasks between Mora and Sora is presented in Table 7. This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake. This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
Tasks Example Sora Mora Others
Text-to-video Generation [Uncaptioned image] (Girdhar et al., 2023; Wang et al., 2023a; Chen et al., 2024a; Ma et al., 2024b)
Image-to-Video Generation [Uncaptioned image] (Blattmann et al., 2023; pik, ; Gen, a)
Extend Generated Videos [Uncaptioned image] -
Video-to-Video Editing [Uncaptioned image] (Molad et al., 2023; Liew et al., 2023; Ceylan et al., 2023)
Connect Videos [Uncaptioned image] (Chen et al., 2023c)
Simulate Digital Worlds [Uncaptioned image] -
Appendix D Discussion
Advantages of Mora.  Mora introduces a groundbreaking multi-agent framework for video generation, advancing the field by enabling a variety of tasks such as text-to-video conversion and digital world simulation. Unlike closed-source counterparts like Sora, Mora’s open framework offers seamless integration of various models, enhancing flexibility and efficiency for diverse applications. As an open-source project, Mora significantly contributes to the AI community by democratizing access to advanced video generation technologies and fostering collaboration and innovation. Future research is encouraged to improve the framework’s efficiency, reduce computational demands, and explore new agent configurations to enhance performance.
Limitations of Mora.  Mora faces significant limitations, including challenges in collecting high-quality video datasets due to copyright restrictions, resulting in difficulties in generating lifelike human movements. Its video quality and length capabilities fall short compared to Sora, with noticeable degradation beyond 12 seconds. Mora also struggles with interpreting and rendering motion dynamics from prompts, lacking control over specific directions. Furthermore, the absence of human labeling information in video datasets leads to results that may not align with human visual preferences, highlighting the need for datasets that adhere to physical laws.
Figure 15: Some video examples generated by Mora.